# Understanding and predicting adsorption energetics on monolayer transition metal dichalcogenides

### v1

@@ Abstract
Two-dimensional (2D) transition metal dichalcogenides (TMDs) have recently been shown to demonstrate non-volatile resistive switching (NVRS), offering significant advantages such as high-density integration and low energy consumption due to their atomic-scale thinness. In this study, we focus on the adsorption and desorption of metal adatoms, which can modulate the electrical resistivity by several orders of magnitude. We develop material-based relationships of the adsorption energy with electronic and atomic structure descriptors by examining the effects of various transition-metal adsorbates on the surface of TMDs. Our results reveal that adsorption energies of transition metals exhibit consistent trends across different TMDs (MoS2, MoSe2, WS2, WSe2) and can be explained using simple descriptors of the atomic and electronic structure. We propose several models to describe this adsorption process, providing a deeper understanding of a crucial step in the resistive switching mechanism based on formation and dissolution of point defects. Finally, we connect our computed adsorption energies to the switching energy. These findings will help guide rational materials selection for the development of NVRS devices using 2D TMDs.
I Introduction
Diversification of memory and storage technologies and improvements of 10$\times$-100$\times$ in capacity and energy efficiency are required, necessitating novel materials platforms [1, 2]. The drive towards smaller and more energy-efficient electronic devices has heightened interest in two-dimensional (2D) materials  [3, 4, 5, 6]. This interest stems from the unique electronic, optical, and scalability properties of the low-dimensional materials [7, 8, 9, 10, 11]. Two-dimensional materials exhibiting NVRS [12, 13, 14] have demonstrated high yields at wafer-scale [15], low cycle-to-cycle and device-to-device variability [16], and fast switching times [17, 18], making them competitive with existing and novel memory and storage technologies and neuromorphic computing applications [19]. As a result, resistive switching based on atomically-thin 2D materials, including the transition metal dichalcogenides (TMDs) [20, 21, 22, 23], are increasingly viewed as promising materials platforms to address the pressing challenges of achieving higher density integration and enhanced energy efficiency in next-generation computing technologies.
@@ I Introduction
How the switching characteristics of memristors depend on the switching and electrode materials is not fully understood. In lateral configurations, where electrodes are placed on the same plane as the TMD layer, Sangwan et al. [24] demonstrated that gate-tunable memristive behavior in monolayer MoS2222 devices allows field-control of switching through electrostatic gating and operates via migration of grain boundaries via bias sweeps in the 10s of volts. For vertical memristors, with electrodes positioned above and below the TMD layer, exhibit distinct out-of-plane switching dynamics. Xu et al. [25] reported vertically stacked memristors using graphene/MoS2/graphene structures, which provided excellent flexibility and endurance and switching voltages in the 100 milliVolt range. Kumar et al. [26] further explored vertically aligned MoS2 layers with metal electrodes, demonstrating that this configuration leads to enhanced on/off ratios and retention times due to improved layer-to-layer coupling. Previous studies have achieved artificial synpases with combining different TMD layers or other 2D materials to form junctions,  [27, 28] but these studies focused on imitating switching characteristics to biological synapses and device integration rather than a materials understanding.
@@
Several mechanisms for resistive switching have also been proposed, including due to the migration of extended defects[24], the generation of vacancies [13], the formation of a conductive filament or bridge [25, 29, 30, 31], the migration of defect ions originating from the electrode [32], and relatedly, the adsorption and desorption of a metal adatom defect on a pre-existing anion vacancy [33, 14, 34, 35, 36]. Thus far, there is a huge variability in reported switching characteristics in low-dimensional materials [5, 37, 38]. For example, MoS2 with gold electrodes experimentally has been reported to have a higher switching ON/OFF ratio compared to WSe2 [14] but the opposite trend is predicted from theory [36], highlighting a critical need to understand the switching and electrode materials and their defect chemistry. Indeed, the plethora of NVRS mechanisms simultaneously highlights the prevalence of resistive switching in low-dimensional materials and the gaps in knowledge of the atomic underpinnings in resistive switching mechanisms at the material level.
@@
Figure 1: (a) Periodic table highlighting the elements studied in this work. Adatoms investigated for adsorption on transition metal dichalcogenides (TMDs) are marked in green. The transition metal (M) and chalcogen (X) elements in the MX2 configuration are highlighted in purple and yellow, respectively. (b) Representative current-voltage (I-V) curve demonstrating resistive switching behavior in a MoS2-based memristor device. The high resistance state (HRS) is indicated in red, corresponding to the system before adatom adsorption, while the low resistance state (LRS) is shown in black, corresponding to the state after adatom adsorption.
@@
In this study, we utilize first-principles calculations based on density functional theory[44, 45] to better understand the defect energetics in the resistive switching mechanism of monolayer transition metal dichalcogenides (TMDs). We focus on resistive switching as mediated by the formation and dissolution of point defects, namely the adsorption and desorption of a (neutral) metal adatom that originates from the electrode material on the surface of the TMD (Figure 1). Such defects are readily expected to form and contribute to resistive switching, as observed computationally [34] and experimentally [46]. In Section III.1, we discuss trends observed in computed adsorption energies for representative monolayer TMDs (MoS2, MoS2, WS2, WSe2) in order to extract meaningful materials descriptors. In order to rationalize trends in adsorption energies, we introduce a model for predicting the adsorption energies of transition metal adsorbates on TMDs, including ones based on the d-band center model inspired by the study of catalysts in Section III.2. We then present a framework for connecting the energetics of adsorption with the switching energy, inspired by the dissociation-diffusion-adsorption model [14] in Section III.4. Finally, we discuss ramifications of the study on considerations for materials selection in novel memory and computing technologies based on resistive switching mediated by point defects (Section IV).
Ultimately, this study seeks to address several key questions: Which adsorbates can induce NVRS? What dictates the adsorption energy and the switching energy? Are there discernible trends among potential materials candidates undergoing switching or for the electrode? By systematically analyzing the adsorption energies and related trends, we aim to accelerate identification of promising materials for future memristive devices based on TMDs or other 2D materials.
II Methods
II.1 Computational Parameters
The calculations in this work were performed using the Vienna Ab initio Simulation Package (VASP.6.3.0) [47, 48, 49] with the Projector Augmented-Wave (PAW) method and the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional [50, 51]. A kinetic energy cut-off of 400 eV was used for the plane-wave basis set expansion. The Brillouin zone was sampled using a 2 $\times$ 2 mesh, and spin polarization was taken into account, as it is expected to be significant when dealing with transition metal adsorbates [52]. Convergence criteria for an energy difference of $10^{-5}$ eV for electronic steps and force difference of 10 meV/Å. For density of states calculations, a $k$-point grid of 4 $\times$ 4 (equivalent to a resolution of 0.07 Å-1) is used, resulting in less than 0.01 eV difference in Fermi energies).
To investigate the overall relationship between the structure and composition of a TMD and adatom complex, we use the adsorption energy defined by Eq. 1. Here, $E_{ads}$ is the adsorption energy, which represents the energy required for the adsorption of a single adatom onto the substrate. The terms $E^{tot}_{ads/sub}$, $E^{tot}_{sub}$, and $E^{tot}_{ads}$ denote the total energies of the structure with the adsorbate, the substrate without the adsorbate, and the chemical potential of the adatom, respectively.
$E_{ads}=E^{tot}_{ads/sub}-E^{tot}_{sub}-E^{tot}_{ads}$ (1)
Figure 2: Relaxed atomic structures of MoS2 with (a) a single sulfur vacancy, (b) gold adsorbate, (c) platinum adsorbate, (d) zirconium adsorbate. The distances between the nearest neighbor molybdenum atoms in each structure are indicated.
In general, the adsorption energy will depend on the chemical potential of the adatom, and hence depend on the bath from where the adatom originates, which is depends on experimental conditions. We explore the effect of the chemical potential ranging from the isolated adatom to that referenced to the bulk metal on the adsorption energy in Section III.1. In Eq. 1, we assume the chemical potential to be effectively that of an isolated adatom in order to focus on the adsorption energetics due to chemical and electronic differences in the adsorbate or monolayer.
To minimize spurious defect-defect interactions, a supercell size of 8 × 8 (192 atoms) was used with a vacuum size of 20 Å. The relaxed structures shown in Figure 2 illustrate the distance between adjacent molybdenum atoms near the defect site. We focus on adsorption of a metal adatom on a pre-existing anion vacancy as the representative point defect complex, as this has been shown in our prior work to be more energetically favorable compared to the adsorption of the adatom on a pristine monolayer [34]. For example, adsorption energies of Ag and Au on pristine MoS2 were calculated to be -0.59 eV and -0.64 eV, respectively, at the site directly above the molybdenum, with similar energies within 0.2 eV at other sites (-0.74 eV above the chalcogen site, -0.73 at the hollow site). In contrast, the adsorption energy at the chalcogen vacancy site was calculated to be -2.56 eV for gold. This is consistent with the dissociation-diffusion-adsorption (DDA) model [14] presented in Section III.4, where the mechanism involves the dissociation of the adatom from the electrode, diffusion across the TMD, and eventual adsorption at an anion vacancy site (see also Section III.4).
II.2 The d-band model
We apply the d-band model to rationalize trends in adsorption energy as they relate to resistive switching applications. The d-band model is based on the Newns-Anderson model [62, 63], which correlates the energy levels of the adsorbate with the electronic structure of the substrate. Subsequent refinements to this model have enhanced its accuracy and applicability, such as the inclusion of spin [52] and the inclusion of structural information via compound descriptors [64], and we utilize aspects of these refinements in our study. This approach utilizes the energy level of the d-band center and the shift induced by the adsorbate as key descriptors for adsorption energy.
@@ II.2 The d-band model
where $\varepsilon$ represents the energy eigenvalues, and $n(\varepsilon)$ is the density of states of the bands with projected d-character. In our calculations, we define the d-band center over a fixed energy range that includes both regions below and above the Fermi level between the different adsorbates. This approach allows for a direct comparison of d-band centers across different adsorbates. By considering a sufficiently broad energy range that spans across the occupied and occupied bands within the band gap, the d-band center reflects the overall electronic environment of the system. Here in this study, we choose an energy range of -10 eV to 1.5 eV (referencing EF = 0 eV). This strategy enables an understanding of how the electronic structure influences adsorption energy.
Here, $f_{\sigma}$ represents the fractional occupancy of spin state $\sigma$, and $\varepsilon_{d\sigma}$ is the conventional Hammer-Nørskov d-band center for spin state $\sigma$. The final term, $\mu$, is the reduced fractional moment, defined as $\mu=\frac{f_{\uparrow}-f_{\downarrow}}{f_{\uparrow}+f_{\downarrow}}$. The first term in the equation gives the spin-averaged d-band center, where the spin-up and spin-down channels are averaged, while the second term introduces a shift in the d-band center that depends on spin polarization.
Apart from the d-band center, other parameters in the d-band model include the orbital occupation, the coupling between the electronic states of the adsorbate and substrate, and the energy associated with the orthogonalization of the states between the adsorbate and substrate. These descriptors are largely based on the electronic structure and thus suggest a generality of the d-band model beyond catalysis, extending to applications such as resistive switching based on the formation of point defects.
As a complement to the d-band center model, we also utilize crystal orbital Hamiltonian population analysis (COHP) to further study the bonding nature of the adsorbate on the various TMDs [66]. These calculations are performed using the Local-Orbital Basis Suite Towards Electronic-Structure Reconstruction (LOBSTER) package [67, 68, 69, 70, 71]. Crystal Orbital Hamiltonian Population (COHP) analysis is a method used to distinguish between bonding, non-bonding, and antibonding interactions between adjacent atoms by analyzing the projected density of states (DOS) weighted by Hamiltonian matrix elements. These elements are related to the overlap integrals between atomic orbitals, allowing for a quantitative assessment of the bonding characteristics in a crystal or molecule. Specifically, the Hamiltonian element corresponds to the interaction strength between pairs of atomic orbitals on neighboring atoms, allowing COHP to quantify the extent of bonding (stabilizing) and anti-bonding (destabilizing) character in a material’s electronic structure. This method provides a detailed understanding of the nature of chemical bonding within the material by linking the DOS with the underlying electronic interactions dictated by the Hamiltonian.
Finally, we use the Sure Independence Screening and Sparsifying Operator (SISSO) [72] to develop interpretable descriptors that leverage chemical intuition to predict adsorption energy (see Table S2 in the SI for the complete list of descriptors). SISSO is a symbolic regression method that systematically combines chemically relevant primary features—such as ionization energy, electronegativity, and atomic radii—using mathematical operations to generate a vast space of potential descriptors. By creating a large pool of candidate features and applying a sparsifying operator, SISSO identifies a minimal set of the most predictive and interpretable descriptors, effectively capturing nonlinear relationships between input features and adsorption energy. This approach ensures the resulting models are physically meaningful and avoids overfitting by selecting descriptors that generalize well across different datasets. The strength of SISSO lies in its ability to reduce the problem’s dimensionality while preserving essential physics, distilling complex material behaviors into a few key descriptors. This enhances predictive accuracy and provides significant insights into the determinants of adsorption energy. Further details on the implementation are provided in Supplementary Information (SI).
III Results
III.1 Trends in the adsorption energy
[…]
@@ III.1 Trends in the adsorption energy
These observed trends, on first examination, are correlated with the atomic electron configuration of the adatom, since one major difference among the adsorbate-substrate pairs in this study is the variation in the valence electron configurations of the adsorbates. When a chalcogen vacancy is present, the cation in the TMD is under-coordinated. The chalcogen vacancy leads to two excess electrons, which alters the electronic environment of the transition metal. When a metal adatom is introduced onto the TMD surface, it can interact with the unsaturated Mo atom, potentially altering its electronic environment further. Metals like Zn, Cd, and Hg, which have fully filled d-orbitals and s-orbitals, exhibit the weakest adsorption energies and, in some cases, even positive adsorption energies. A positive adsorption energy in this context indicates that these metals do not readily adsorb onto the vacancy. Due to their stable electron configurations, these metals do not readily donate or accept electrons, resulting in minimal interaction with the unsaturated Mo atom. Similarly, metals such as Ag, Au, and Cu, which have nearly filled d-orbitals, also show relatively weak adsorption energies. Although Cu is not a noble metal, its nearly filled d-orbital leads to weaker interactions with the TMD surface, similar to the behavior of Ag and Au. In contrast, metals with partially filled d-orbitals, such as Fe, typically exhibit intermediate adsorption energies. The overall trend of stronger adsorption energies continues when examining the early transition metals, which have mostly unfilled d-orbitals. Thus, there is a correlation with the extent of electron filling of the adatom mediating the adsorption energy.
Figure 3: Adsorption energies (left, blue axis) of various transition metal adsorbates on a single sulfur vacancy in MoS2. Adsorbates in each plot are sorted and colored by their periodic group. The right axis (red circles) displays the Bader charge gain or loss of the adsorbate relative to its neutral atomic state, with negative values indicating charge gain and positive values indicating charge loss.
@@
We note that the adsorption energy will vary depending on the value of the chemical potential, which will depend on experimental and device conditions. Thus, we look at two limits of the chemical potential where the adatom is referenced to an isolated adatom and to the monatomic metal bulk. For the adsorption energies in Figure 3, the chemical potential is chosen as the isolated adatom, assuming dissociation of the adatom from the electrode has already occurred. When instead using the chemical potential referenced to the monatomic metal bulk, the adsorption energies can vary up to 10.93 eV, as seen in the case of tungsten (W), which had one of the largest differences between the two reference states. This difference highlights how the choice of chemical potential can affect the strength of the adsorbate-substrate interaction. A plot showing the variation of the adsorption energy with the chemical potential referenced to that of the bulk metal is provided in Figure S5 of the SI for MoS2, MoSe2, WS2, and WSe2. Referencing the chemical potential to the bulk metal results in more positive (i.e., less exothermic) adsorption energies, and reflects a scenario where the metal adatom is part of a bulk reservoir. To focus on the differences in adsorption energy arising from the interactions between the monolayer and adsorbate, we continue our discussion using the chemical potential of the isolated adatom.
III.2 Rationalizing trends in adsorption energy

### v2

@@ Abstract
Two-dimensional (2D) transition metal dichalcogenides (TMDs) have emerged as promising candidates for non-volatile resistive switching (NVRS) due to their atomic-scale thickness, enabling high-density integration and low energy consumption. This study explores how metal adatom adsorption and desorption modulate resistivity in these materials. By examining a range of transition-metal adsorbates on MoS2, MoSe2, WS2, and WSe2, we uncover material-specific trends in adsorption energies that correlate with electronic and atomic structure descriptors. These trends are captured by simple models that elucidate how adsorption influences the formation and dissolution of point defects, a key mechanism in NVRS. Our results highlight consistent periodic behavior across TMDs and demonstrate that adsorption energy is a useful predictor of switching energy. This understanding provides insight into atomic-scale switching phenomena and offers design principles for selecting TMD-adsorbate pairs to optimize NVRS device performance.
This is the accepted manuscript of the article published in npj 2D Materials and Applications (2025), DOI: 10.1038/s41699-025-00579-9.
1 Introduction
@@
Diversification of memory and storage technologies and improvements of 10$\times$-100$\times$ in capacity and energy efficiency are required, necessitating novel materials platforms [1, 2]. The drive towards smaller and more energy-efficient electronic devices has heightened interest in two-dimensional (2D) materials  [3, 4, 5, 6]. This interest stems from the unique electronic, optical, and scalability properties of the low-dimensional materials [7, 8, 9, 10, 11]. Two-dimensional materials exhibiting non-volatile resistive switching (NVRS) [12, 13, 14] have demonstrated high yields at wafer-scale [15], low cycle-to-cycle and device-to-device variability [16], and fast switching times [17, 18], making them competitive with existing and novel memory and storage technologies and neuromorphic computing applications [19]. As a result, resistive switching based on atomically-thin 2D materials, including the transition metal dichalcogenides (TMDs) [20, 21, 22, 23], are increasingly viewed as promising materials platforms to address the pressing challenges of achieving higher density integration and enhanced energy efficiency in next-generation computing technologies. How the switching characteristics of memristors depend on the switching and electrode materials is not fully understood. In lateral configurations, where electrodes are placed on the same plane as the TMD layer, Sangwan et al. [24] demonstrated that gate-tunable memristive behavior in monolayer MoS2 devices allows field-control of switching through electrostatic gating and operates via migration of grain boundaries via bias sweeps in the tens of volts. Vertical memristors with electrodes positioned above and below the TMD layer exhibit distinct out-of-plane switching dynamics. Xu et al. [25] reported vertically stacked memristors using graphene/MoS2/graphene structures, which provided excellent flexibility and endurance and switching voltages in the 100 milliVolt range. Kumar et al. [26] further explored vertically aligned MoS2 layers with metal electrodes, demonstrating that this configuration leads to enhanced on/off ratios and retention times due to improved layer-to-layer coupling. Previous studies have achieved artificial synapses with combining different TMD layers or other 2D materials to form junctions  [27, 28], but these studies focused on imitating switching characteristics of biological synapses and device integration rather than a materials understanding.
@@
Several mechanisms for resistive switching have also been proposed, including the migration of extended defects[24], the generation of vacancies [13], the formation of a conductive filament or bridge [25, 29, 30, 31], the migration of defect ions originating from the electrode [32], and relatedly, the adsorption and desorption of a metal adatom defect on a pre-existing chalcogen vacancy [33, 14, 34, 35, 36]. Thus far, there is a huge variability in reported switching characteristics in low-dimensional materials [5, 37, 38]. For example, MoS2 with gold electrodes has been reported to have a higher switching ON/OFF ratio compared to WSe2 in experiment [14] but the opposite trend is predicted from theory [36], highlighting a critical need to understand the switching and electrode materials and their defect chemistry. Indeed, the plethora of NVRS mechanisms simultaneously highlights the prevalence of resistive switching in low-dimensional materials and the gaps in knowledge of the atomic underpinnings in resistive switching mechanisms at the material level.
@@
In this study, we utilize first-principles calculations based on density functional theory[44, 45] to better understand the defect energetics in the resistive switching mechanism of monolayer transition metal dichalcogenides (TMDs). We focus on resistive switching as mediated by the formation and dissolution of point defects, namely the adsorption and desorption of a (neutral) metal adatom that originates from the electrode material on the surface of the TMD (Figure 1). Such defects are readily expected to form and contribute to resistive switching, as observed computationally [34] and experimentally [46]. In Section 2.1, we discuss trends observed in computed adsorption energies for representative monolayer TMDs (MoS2, MoSe2, WS2, WSe2) in order to extract meaningful materials descriptors. In order to rationalize trends in adsorption energies, we introduce a model for predicting the adsorption energies of transition metal adsorbates on TMDs, including ones based on the d-band center model inspired by the study of catalysts in Section 2.2. We then present a framework for connecting the energetics of adsorption with the switching energy, defined as the energy required to change the resistive state of the device, inspired by the dissociation-diffusion-adsorption model [14] in Section 2.4. Finally, we discuss ramifications of the study on considerations for materials selection in novel memory and computing technologies based on resistive switching mediated by point defects (Section 3).
Ultimately, this study seeks to address several key questions: Which adsorbates can induce NVRS? What dictates the adsorption energy and the switching energy? Are there discernible trends among potential materials candidates undergoing switching or for the electrode? Our study reveals that there are broad, underlying material trends in the defect energetics, which are correlated with the resistive switching behavior of the device. We utilize theoretical models such as the d-band center theory, charge analysis, chemical bonding analysis, and compressed sensing techniques to rationalize trends in adsorption energetics. This study demonstrates that physical models of adsorption can be used to understand trends in energetics related to the switching energy in neuromorphic devices. Notably, unlike conventional applications of the d-band model involving organic adsorbates, we demonstrate that such adsorption physics can also apply to transition metal adsorbates. Our approach simplifies the modeling of adsorption energy from existing models [43], which will provide a valuable tool for materials-driven design. By addressing these questions, this study aims to accelerate the identification of materials for memristive devices based on TMDs and other 2D materials.
2 Results
2.1 Trends in the adsorption energy
We examine the adsorption energies of different transition metal adatoms on TMD monolayers, revealing the underlying principles of their interactions. In particular, we focus on the adsorption of transition metal adatoms onto chalcogen vacancies, an intrinsic defect present in MoS2, MoSe2, WS2, and WSe2 samples [47, 34]. Our computed adsorption energies align well with prior computational studies. For instance, the adsorption energy of gold onto a sulfur vacancy is found to be -2.64 eV, which is within $\pm$0.1 eV of values reported in the literature with slight variations arising from factors such as k-point grid, supercell size, and choice of chemical potential [48, 43, 14].
@@ 2.1 Trends in the adsorption energy
As shown in Figure 2, adsorption energy and Bader charge gain or loss are loosely inversely correlated. Transition metal adsorbates that exhibit greater charge loss generally show stronger adsorption energies. The earliest transition metals (Sc, Y) exhibit weaker adsorption energies than anticipated ($E^{\text{Sc}}_{ads}=-5.57$ eV, $E^{\text{Y}}_{ads}=-5.83$ eV) compared to later transition metal adatoms that lose less charge but have stronger adsorption interactions (e.g., $E^{\text{Ti}}_{ads}=-6.13$ eV, $E^{\text{Zr}}_{ads}=-7.66$ eV, $E^{\text{Hf}}_{ads}=-7.31$ eV). However, the charge state of the adsorbate is only one aspect of the overall interaction. Notable exceptions include osmium (Os), which exhibits minimal charge gain but still has a strong adsorption energy.
When a chalcogen vacancy is present, the cation in the TMD becomes under-coordinated, leaving two excess electrons that alter the electronic environment of the transition metal adsorbate. This altered environment mediates the interaction strength between the adsorbate and the substrate, which depends not only on charge loss or gain but also on the adsorbate’s ability to hybridize effectively with the substrate.
@@
These trends in adsorption energy can be partially explained by the variation in valence electron configurations among adsorbates. Metals like Zn, Cd, and Hg, which have fully filled d-orbitals and s-orbitals, exhibit the weakest adsorption energies and, in some cases, even positive adsorption energies. A positive adsorption energy in this context indicates that these metals do not readily adsorb onto the vacancy. Their stable electron configurations result in lessened interaction with the unsaturated Mo atom, as these metals do not readily donate or accept electrons. Similarly, metals such as Ag, Au, and Cu, which have nearly filled d-orbitals, also show relatively weak adsorption energies. Although Cu is not a noble metal, its nearly filled d-orbital leads to weaker interactions with the TMD surface, similar to the behavior of Ag and Au.
In contrast, metals with partially filled d-orbitals, such as Fe, typically exhibit intermediate adsorption energies. The overall trend of stronger adsorption energies continues when examining the early transition metals, which have mostly unfilled d-orbitals. There is also an apparent periodicity in the adsorption energies with the extent of Bader charge loss or gain within each group. The periodicity in adsorption energy is altered when charge loss transitions into charge gain, which is notable for adsorbates such as Os, Ir, Pt, and Au. As discussed in Section 2.2, multiple factors play a role in the adsorption energy, such as electron filling, orbital hybridization, and orbital geometry.
@@
We note that the adsorption energy will vary depending on the value of the chemical potential, which will depend on experimental and device conditions. Thus, we look at two limits of the chemical potential where the adatom is referenced to an isolated adatom and to the monatomic metal bulk. For the adsorption energies in Figure 2, the chemical potential is chosen as the isolated adatom, assuming dissociation of the adatom from the electrode has already occurred. When instead using the chemical potential referenced to the monatomic metal bulk, the adsorption energies can vary up to 10.93 eV, as seen in the case of tungsten (W), which had one of the largest differences between the two reference states. This difference highlights how the choice of chemical potential can affect the strength of the adsorbate-substrate interaction. A plot showing the variation of the adsorption energy with the chemical potential referenced to that of the bulk metal is provided in Supplementary Figure 7 for MoS2, MoSe2, WS2, and WSe2. Referencing the chemical potential to the bulk metal results in more positive (i.e., less exothermic) adsorption energies, and reflects a scenario where the metal adatom is part of a bulk reservoir. To focus on the differences in adsorption energy arising from the interactions between the monolayer and adsorbate, we continue our discussion using the chemical potential of the isolated adatom.
2.2 Rationalizing trends in adsorption energy
To rationalize the observed trends, we employ the d-band center model, which combines the effects of electron filling and the formation of bonding and anti-bonding orbitals between the adsorbate and the substrate (also see Section 4.2).